One kg of a diatomic gas is at a pressure of 8 × 10(4) N/m2. The density of the gas is 4 kg/m(3). What is the energy of the gas due to its thermal motion ?

Grade:11

2 Answers

Arun
3 years ago
For a diatomic molecular gas, internal energy per molecule is
U=25NKBT
PV=NKBT
From equation is (i) and (ii)
U=25PV
Here, P=8×104N m2,ρ=4kg m3m=1kg
U=25×8×104×41=5×104J
Vikas TU
3 years ago
Internal energy due to thermal motion is equal to 5/2 × P × V
Mass = 1kg
Density = 4kg / m^3
Volume = 1/5 m^3
Pressure = 8 × 10^4 n/m^2
Energy = 5/2 × 1/5 × 8 × 10^4
Answer =Energy = 4 × 10^4 Joules